At a discount of 4% for 90-day treasury bill, an investor paid 970.10 pesos. How much will the investor received at maturity??
1
Expert's answer
2020-11-11T18:43:43-0500

Y= (970.10/x) (360/90) =0.04


Yd= (970.10/x)(4) =0.04


Y= 3880.4/x = 0.04


3880.4/x = 0.04


0.04x = 3880.10


X = 3880.10/0.04


X = 97010 pesos